OpenStudy (anonymous):

Find the equation of the line that contains the point (20,26) and is parallel to the axis. Write the equation in the form y+mx+b and identify m and b.

OpenStudy (campbell_st):

which axis is the line parallel to..?

OpenStudy (anonymous):

The x axis, sorry!

OpenStudy (campbell_st):

the x axis has a slope m = 0... to the line is y = 26  you can use the slope intercept form with x = 20, y = 26 and m = 0 26 = 0 * 20 + b b = 26 then the line is y = 0x + 26 or y = 26